# Emotionally Numb or Empathetic? Evaluating How LLMs Feel Using EmotionBench

### v5

@@ 7 Conclusion
Appendix A More Information on Psychology
@@ Appendix E Prompting LLMs To Be Emotionally Stable
Appendix F Ethics Statement and Broader Impacts
F.1 Safeguards on Human Subjects
This study involves a survey requiring human subjects to imagine being in situations that could elicit negative emotions such as anger, anxiety, and fear. This process introduces a few ethical concerns. First, this process could hurt the mental health of human subjects. To alleviate the possibility, we take the following actions: (1) We require subjects to be free of any ongoing mental illness. (2) We inform subjects about the nature of the survey in advance, including the potential risks of emotional distress. (3) We allow all subjects to quit at any time. (4) We provide mental support and let subjects report any illness after the survey. Fortunately, no subjects reported such kind of mental illness. Another concern is related to the privacy issue during the collection of data. Our questionnaire is entirely anonymous to safeguard subjects’ privacy and confidentiality.
F.2 Impacts on LLM Developers and Users

### v6

@@ 7 Conclusion
Acknowledgments
We would like to express our sincere gratitude to Xiaoyuan Liu and Pinjia He from the Chinese University of Hong Kong, Shenzhen, for their valuable assistance during the rebuttal process. The paper is supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (No. CUHK 14206921 of the General Research Fund).
Appendix A More Background from Psychology
@@ Appendix E Prompting LLMs To Be Emotionally Stable
Appendix F Tuning LLMs To Align with Humans
We conduct an experiment using the GPT-3.5-Turbo model and the LLaMA-3.1-8B model. Our EmotionBench (1,266 human responses) is split into 866 samples for fine-tuning and 400 for testing. The following hyperparameters are used: n_epochs$\ =3$, batch_size$\ =1$, and learning_rate_multiplier$\ =2$ for GPT-3.5-Turbo, and learning_rate$\ =5\times 10^{-5}$, per_device_train_batch_size$\ =2$, and num_train_epochs$\ =3$ for LLaMA-3.1-8B. For LLaMA-3.1, we apply the Low-Rank Adaptation (LoRA) Hu et al. (2022) technique. Table 15 compares the performance of the vanilla and fine-tuned models against human baseline, specifically in terms of negative affect scores from the test set.
Table 15: Performance comparison of vanilla (marked as V) and fine-tuned (marked as FT) GPT-3.5 and LLaMA-3.1 models on negative affect scores.
Models Human GPT-3.5 (V) GPT-3.5 (FT) LLaMA-3.1-8B (V) LLaMA-3.1-8B (FT)
Default (N) $14.2_{\pm 6.4}$ $25.9_{\pm 0.3}$ $10.6_{\pm 0.5}$ $33.0_{\pm 4.5}$ $10.3_{\pm 1.1}$
Evoked (N) $25.9_{\pm 9.7}$ $24.8_{\pm 8.5}$ $25.2_{\pm 9.6}$ $36.5_{\pm 7.7}$ $15.0_{\pm 6.4}$
The results show that fine-tuned models align more closely with human emotional responses in both default and emotion-evoked states. Notably, fine-tuning the models using our dataset significantly improved emotional alignment, particularly for the LLaMA-3.1 model, which reduced its negative affect score from $33.0$ to $10.3$ in the default state. These findings demonstrate the effectiveness of our EmotionBench dataset in enhancing models’ emotional alignment with human norms.
Appendix G Ethics Statement and Broader Impacts
G.1 Safeguards on Human Subjects
This study involves a survey requiring human subjects to imagine being in situations that could elicit negative emotions such as anger, anxiety, and fear. This process introduces a few ethical concerns. First, this process could hurt the mental health of human subjects. To alleviate the possibility, we take the following actions: (1) We require subjects to be free of any ongoing mental illness. (2) We inform subjects about the nature of the survey in advance, including the potential risks of emotional distress. (3) We allow all subjects to quit at any time. (4) We provide mental support and let subjects report any illness after the survey. Fortunately, no subjects reported such kind of mental illness. Another concern is related to the privacy issue during the collection of data. Our questionnaire is entirely anonymous to safeguard subjects’ privacy and confidentiality. The Survey and Behavioural Research Ethics (SBRE) Committee from the Chinese University of Hong Kong has granted approval for this study, titled “Exploring Human Emotional Responses to Diverse Situations,” with the reference number of SBRE‐23‐0696.
G.2 Impacts on LLM Developers and Users